# Mind the Gap: A Review of Arabic Post-Training Datasets and Their Limitations

### v1

@@ 5 Case Study: Safety and Cultural Alignment
This underrepresentation is especially concerning given the importance of cultural sensitivity and safety in Arabic-speaking contexts, where linguistic, societal, and religious norms differ greatly from dominant English-based benchmarks. The lack of culturally aware and safety-focused datasets increases the risk of deploying misaligned or even harmful NLP systems. To address these blind spots, we recommend prioritizing the development of high-quality datasets tailored to Arabic cultural contexts and safety concerns, ensuring that future models are not only technically robust but also ethically and socially aligned.
@@ 6 Recommendations and Future Directions
The findings of this survey highlight the strategic importance of post-training datasets for advancing Arabic-centric LLMs. While the existing resources on Hugging Face provide a starting point, they fall short in coverage, documentation quality, cultural alignment, and scientific rigor. To address these limitations and accelerate the development of Arabic LLMs, we offer the following forward-looking recommendations, structured around priority domains, practical dataset creation strategies, and principles for collaborative research.
@@ 6.1 High-Priority Domains for Future Post-Training Datasets
This subsection outlines specific domains in Arabic post-training that are currently underrepresented or entirely missing, yet are crucial for building capable, safe, and culturally aligned Arabic LLMs. These domains should be prioritized in future dataset development initiatives due to their strategic importance and lack of coverage.
@@
  • Summarization: While moderately covered, many existing datasets lack consistency in documentation, linguistic variety, and practical relevance to real-world summarization use cases.
  • Cultural Alignment: Data that reflects nuanced Arab world values, norms, and social constructs is crucial for building culturally sensitive systems.
  • Dialog/Conversation: This domain suffers from very limited coverage and low-quality documentation. Rich, Dialect-sensitive dialogue datasets are essential for improving conversational fluency and natural interaction in Arabic LLMs.
@@
  • Robustness & Safety: Despite its importance for responsible AI development, the availability of high-quality Arabic datasets in this domain remains limited.
  • Function Calling: Essential for tool-augmented NLP and API-connected LLMs, yet currently nonexistent in Arabic post-training resources.
  • Ethics, Bias, and Fairness: Arabic datasets in this area are extremely limited, despite growing ethical concerns in global LLM deployment.
  • Code Generation: There are currently no Arabic post-training datasets supporting code generation.
  • Official Documentation: This domain is completely absent from current post-training resources, though critical for building LLMs that can handle structured instructions, policies, or formal content.
@@ 6.2 Practical Guidelines for Building Arabic Post-Training Datasets
This subsection focuses on practical and scalable methods for creating Arabic post-training datasets. These guidelines are intended for researchers and developers who aim to build new resources and address domain-specific gaps identified in Section 7.1. The listed methods are grounded in existing tools, community collaboration, and modern data generation strategies.
@@
A crowdsourced annotation platform can empower native speakers to label data along cultural and contextual dimensions. By providing well-defined annotation guidelines, especially on culturally sensitive topics, the platform can produce high-quality datasets with rich socio-cultural nuance.
@@ 6.3 Recommendations for Future Research and Collaboration
This final subsection presents high-level, strategic guidance for the broader research community. These recommendations emphasize principles such as authenticity, cultural representation, and open collaboration. They are intended to shape future initiatives and encourage ethical, inclusive, and sustainable development of Arabic post-training datasets.
  • Prioritize Missing Domains: Direct funding, research, and community efforts toward domains with little to no coverage in Arabic (e.g., function calling, code generation).
  • Promote Authenticity over Translation: Native Arabic content should be favored to avoid loss of context, nuance, or cultural misalignment present in translated material.
  • Incorporate Cultural Context: Datasets should reflect the ethical, religious, and societal values of the Arab world to ensure cultural robustness in AI outputs.
  • Broaden Linguistic Representation: Both MSA and regional dialects should be represented in future dataset development to support real-world use cases across the Arab region.
  • Foster Open Collaboration and Transparency: Dataset creators are encouraged to share licensing details, evaluation metrics, and use-case documentation to increase reproducibility and adoption.
@@ 7 Conclusion
In this paper, we conducted the first systematic survey of publicly available Arabic post-training datasets hosted on the Hugging Face Hub, with a focus on evaluating their quality, coverage, licensing transparency, and scientific contribution across twelve key LLM capabilities. Our findings reveal several critical gaps, most notably the near absence of datasets in high-impact domains such as function calling, code generation, ethical alignment, and official documentation.
Despite the growing importance of post-training in aligning LLMs with human intent, Arabic remains substantially underrepresented in this space. Many existing datasets suffer from limited documentation, outdated maintenance, and low practical adoption. These shortcomings hinder the advancement of robust, culturally aligned, and ethically grounded Arabic LLMs.
We proposed a set of high-priority domains that require urgent dataset development and provided practical, scalable guidelines for building Arabic post-training resources through community collaboration, hybrid human–LLM annotation, and synthetic data generation. Additionally, we outlined strategic recommendations for promoting native content, cultural awareness, and linguistic diversity in future dataset creation efforts.
@@
While this study provides the first structured review of Arabic post-training datasets, it is subject to several limitations. First, this review covers only datasets openly available on Hugging Face, omitting any private or gated resources.
Second, our evaluation relies heavily on metadata and README documentation, which may not always accurately reflect the actual quality or usability of the datasets. Some datasets may be under-documented despite being high quality in practice, and others may appear polished but lack effective downstream utility.
@@ Appendix A Evaluation Criteria
Table 4: Scoring rubric for evaluating Arabic datasets based on popularity, adoption, recency and maintenance, licensing transparency, and scientific contribution. Each criterion is scored individually and mapped to a qualitative level (High, Medium, or Low).
Evaluation Criteria Score Total Score Level Popularity Dataset’s Number of Likes Number of Likes $200\leq\mathrm{Score}$ $200100\leq\mathrm{Score}<$ 200 $\mathrm{Score}<100$ High Medium Low Dataset’s Number of Downloads Number of Downloads Adoption Number of Used Models Number of Models $50\leq\mathrm{Score}$ $5020\leq\mathrm{Score}<$ 50 $\mathrm{Score}<20$ High Medium Low Number of Used Spaces Number of Spaces Recency & Maintenance Dataset’s Last Modified Date Last Modified – Collection Date $\mathrm{Score}\leq 6\,$Mo 666\,6Mo $<\mathrm{Score}\leq 12\,$Mo $\mathrm{Score}>12\,$Mo High Medium Low Licensing Transparency Dataset card states the license License Name Known license ’unknown’/’other’ ’none’ High Medium Low Metadata tags state the license License Name Scientific Contribution Dataset card includes ACL Papers ACL Papers $3\leq\mathrm{Score}$ $31\leq\mathrm{Score}<$ 3 $\mathrm{Score}=0$ High Medium Low Metadata tags include ArXiv Papers ArXiv Papers Metadata tags include a DOI Object DOI Object

### v2

@@ 5 Case Study: Safety and Cultural Alignment
This underrepresentation is especially concerning given the importance of cultural sensitivity and safety in Arabic-speaking contexts, where linguistic, societal, and religious norms differ greatly from dominant English-based benchmarks. The lack of culturally aware and safety-focused datasets increases the risk of deploying misaligned or even harmful NLP systems, like LLMs. To address these blind spots, we strongly recommend prioritizing the development of high-quality datasets tailored to Arabic cultural contexts and safety concerns, ensuring that future models are not only technically robust but also ethically and socially aligned.
@@ 6 Recommendations and Future Directions
The findings of this review highlight the strategic importance of post-training datasets for advancing Arabic-centric LLMs. While the existing resources on Hugging Face Hub provide a starting point, they fall short in coverage, documentation quality, cultural alignment, and scientific rigor. To address these limitations and accelerate the development of Arabic LLMs, we offer the following forward-looking recommendations, structured around priority domains, practical dataset creation strategies, and principles for collaborative research.
@@ 6.1 High-Priority Domains for Future Post-Training Datasets
This subsection outlines specific domains in Arabic post-training that are currently underrepresented or entirely missing, yet are crucial for building capable, safe, and culturally aligned Arabic LLMs. These domains should be prioritized in future post-training dataset development initiatives due to their strategic importance and lack of coverage.
@@
  • Summarization: While moderately covered, many existing datasets lack consistency in documentation, linguistic variety, and practical relevance to real-world use cases.
  • Cultural Alignment: Data that reflects nuanced Arab world values, norms, and social constructs is crucial for building culturally sensitive NLP systems and applications.
  • Dialog/Conversation: This domain suffers from very limited coverage and low-quality documentation and annotation. Rich, dialect-sensitive dialogue datasets are essential for improving conversational fluency and natural interaction in Arabic-centric LLMs.
@@
  • Robustness & Safety: Despite its importance for responsible AI development, the availability of high-quality Arabic post-training datasets in this domain remains limited.
  • Function Calling: Essential for tool-augmented NLP and API-connected LLMs, yet currently nonexistent in public Arabic post-training resources.
  • Ethics, Bias, and Fairness: Arabic datasets in this area are extremely limited, despite growing ethical concerns in global LLM adoption, development, and deployment.
  • Code Generation: There are currently no open Arabic datasets supporting code generation.
  • Official Documentation: This domain is completely absent from current post-training resources, although critical for building capable LLMs that can handle policies, manuals, formal content, or structured instructions.
@@ 6.2 Practical Guidelines for Building Arabic Post-Training Datasets
This subsection focuses on practical and scalable methods for creating Arabic post-training datasets. These guidelines are intended for researchers and developers, who aim to build new resources and address domain-specific gaps. The listed methods are grounded in existing tools, community collaboration, and modern data generation strategies.
@@
A crowd-sourced annotation platform can empower native speakers to label data along cultural and contextual dimensions. By providing well-defined annotation guidelines, especially on culturally sensitive topics, the platform can produce high-quality datasets with rich sociocultural nuance.
@@ 6.3 Recommendations for Future Research and Collaboration
This final subsection presents high-level, strategic guidance for the broader research community. These recommendations emphasize principles like authenticity, cultural representation, and open collaboration. They are intended to shape future initiatives and encourage ethical, inclusive, and sustainable development of Arabic post-training datasets.
  • Prioritize Missing Domains: Direct funding, research, and community efforts toward domains with little to no coverage in Arabic (e.g., Function Calling and Code Generation).
  • Promote Authenticity over Translation: Native Arabic content should be favored to avoid loss of context, nuance, or cultural misalignment present in translated material. While translated datasets can serve as a temporary bridge to address data scarcity, they fundamentally compromise the linguistic and cultural integrity essential for powerful Arabic LLMs. Native Arabic content preserves cultural subtleties, idiomatic expressions, and the language’s unique morphological complexity that translation inevitably distorts. In culturally sensitive domains—including religious discourse, legal frameworks, and social interactions—native content ensures terminological accuracy and cultural appropriateness that directly impacts model performance and user acceptance. Thus, we recommend prioritizing investment in native Arabic dataset creation as a sustainable strategy for developing LLMs that authentically serve Arabic-speaking communities rather than imposing linguistic patterns from other language contexts.
  • Incorporate Cultural Context: Datasets should reflect ethical, religious, and societal views, values, and cultures of the Arab world to ensure cultural robustness in AI outputs.
  • Broaden Linguistic Representation: Both Modern Standard Arabic (MSA) and regional Dialectal Arabic (DA) should be represented in future dataset development to support real-world use cases across the Arab region.
  • Foster Open Collaboration and Transparency: Dataset creators are encouraged to share licensing details, evaluation metrics, and use-case documentation to increase reproducibility, transparency, and adoption.
  • Investigate Dataset-Performance Relationships: Future research should investigate relationships between our categorized dataset characteristics and actual model performance. Such studies could leverage our framework to conduct controlled experiments across task categories, establishing empirical relationships between dataset quality metrics and model effectiveness. This would provide valuable guidance for dataset creators and model developers in the Arabic NLP community.
@@ 7 Conclusion
In this paper, we conducted the first systematic survey of publicly available Arabic post-training datasets hosted on the Hugging Face Hub, with a focus on evaluating their quality, coverage, licensing transparency, and scientific contribution, across 12 key LLM capabilities. Our findings reveal several critical gaps, most notably the near absence of datasets in high-impact domains, such as Function Calling, Code Generation, Ethical Alignment, and Official Documentation. Despite the growing importance of post-training in aligning LLMs with human intent, Arabic remains substantially underrepresented in this space. Many existing datasets suffer from limited documentation, outdated maintenance, and low practical adoption. These shortcomings hinder the advancement of robust, culturally aligned, and ethically grounded Arabic LLMs.
We proposed a set of high-priority domains that require urgent dataset development and provided practical, scalable guidelines for building Arabic post-training resources through community collaboration, hybrid human–LLM annotation, and synthetic data generation. Additionally, we outlined strategic recommendations for promoting native content, cultural awareness, and linguistic diversity in future dataset creation efforts. Lastly, we released two open-source demo versions of our dataset collection and evaluation tools to the Arabic NLP research community. The introduction of these tools will facilitate standardized evaluation practices as well as reproducible research. In the near future, we aim to publicly share production versions with detailed documentation to ensure broad accessibility and adoption across research institutions.
@@
While this study provides the first structured review of Arabic post-training datasets, it is subject to several limitations. First, this review covers only datasets openly available on Hugging Face Hub, omitting any private or gated resources.
Second, our collection and evaluation rely heavily on metadata and Dataset Cards (README) documentation, which may not always accurately reflect the actual quality or usability of the datasets. Some datasets may be underdocumented despite being high-quality in practice, and others may appear polished but lack effective downstream utility.
Third, this study does not assess how the reviewed datasets directly impact model performance. While our review provides essential infrastructure for dataset discovery, examining correlations between dataset characteristics and model effectiveness would require extensive computational resources and standardized benchmarking protocols beyond this study’s scope. As such, the current study did not examine the relationship between the reviewed datasets and model performance.
@@ Appendix A Evaluation Criteria
Table 4: Scoring rubric for evaluating Arabic datasets based on popularity, adoption, recency and maintenance, licensing transparency, and scientific contribution. Each criterion is scored individually and mapped to a qualitative level (High, Medium, or Low). The documentation criteria and scoring rubric are previously displayed in Table 1.
Evaluation Criteria Score Total Score Level Popularity Dataset’s Number of Likes Number of Likes $200\leq\mathrm{Score}$ $100\leq\mathrm{Score}<200$ $\mathrm{Score}<100$ High Medium Low Dataset’s Number of Downloads Number of Downloads Adoption Number of Used Models Number of Models $50\leq\mathrm{Score}$ $20\leq\mathrm{Score}<50$ $\mathrm{Score}<20$ High Medium Low Number of Used Spaces Number of Spaces Recency & Maintenance Dataset’s Last Modified Date Last Modified – Collection Date $\mathrm{Score}\leq 6\,$Mo 66\,Mo $\mathrm{Score}\leq 12\,$Mo $\mathrm{Score}>12\,$Mo High Medium Low Licensing Transparency Dataset card states the license License Name Known license ’unknown’/’other’ ’none’ High Medium Low Metadata tags state the license License Name Scientific Contribution Dataset card includes ACL Papers ACL Papers $3\leq\mathrm{Score}$ $1\leq\mathrm{Score}<3$ $\mathrm{Score}=0$ High Medium Low Metadata tags include ArXiv Papers ArXiv Papers Metadata tags include a DOI Object DOI Object